OpenStudy (anonymous):

Give medal and fan Karina is creating two pyramids with the nets below using construction paper. In both nets, all of the triangular faces of the pyramids have a height of 9 inches and a base of 6 inches. Compare the amount of construction paper needed to make each pyramid. The triangular pyramid needs ________square inches of construction paper to make. The square pyramid needs _____ square inches of construction paper to make. The pyramid will need ________ square inches more construction paper to make than the other pyramid.

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@Unanimoose

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Help please

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

@Bee_see

OpenStudy (anonymous):

@pooja195

OpenStudy (mathmath333):

\(\large \color{black}{ \normalsize \text{for triangular pyramid , paper needed }\hspace{.33em}\\~\\ =4(A(\triangle)\hspace{.33em}\\~\\ \normalsize \text{and }\hspace{.33em}\\~\\ A(\triangle) =\frac{1}{2}\times\normalsize \text{base}\times \normalsize \text{height}\hspace{.33em}\\~\\ =\frac{1}{2}\times6\times 9\hspace{.33em}\\~\\ }\)

OpenStudy (anonymous):

what?